A plane took 4 h to fly 2200 km from saskatoon to toronto with a tail wind. Thereturn trip, with a head wind, took 5 h. Find the speed of the plane in still air and the wind speed.
Answer by dabanfield@earthlink.net(8)

A plane took 4 h to fly 2200 km from saskatoon to toronto with a tail wind. Thereturn trip, with a head wind, took 5 h. Find the speed of the plane in still air and the wind speed.
Let w be the wind speed and s be the speed of the plane in still air.
Using the formula distance = rate*time, for the trip with the wind we have:
2200 = (s+w)* 4 or
1.) 2200 = 4*s + 4*w
For the trip against the wind we have:
2200 = (s-w) *5 or
2.) 2200 = 5*s - 5*w
Multiplyng 1.) by 5 and 2.) by 4 we have:
3.) 11000 = 20*s + 20w
4.) 8800 = 20*s - 20*w
Adding the respective sides of 3 and 4 we have:
19800 = 40*s
s = 495
Substitute 495 for s in either 1.) or 2.) above and solve for w.